# Spin-operator form factors of the critical Ising chain and their finite volume scaling limits

### v3

@@ 1.3 Square-root products and form-factors at finite NN
For the convenience of readers, below we provide a result list for all the computed form-factors. The results at a generic $N$ are given by Eq. (79), and their scaling limits are in Eq. (107), Eq. (108). It turns out that certain square-root products and their relations are crucial in all the stages of the derivation and appear also in the final results, so we start by introducing all the required square-root products. In our algorithm, they will be naturally generated after inverting certain Cauchy matrices and performing further summations. For $\beta\in S_{+}$ one needs two products
@@
They will be used latter throughout the derivation, in particular, to show a crucial matrix is invertible.
On the other hand, for $\beta\neq 1$, no such simple formulas can be derived for the $p$ function. From the perspective of the $S^{2N}_{\pm}$ circles, this is because all the products are restricted to the upper half-plane. But for both $S_{+}$ and $S_{-}$, the $p$ function can still be simplified into a single integral representation. What one need are the following
@@
where $\cal Z$} can be both $\beta$ and $z$. In the expression, the logarithm is define with the principal branch. The function $\hat{\Gamma}_{N}(\beta)$ and $\hat{\Gamma}_{N}(z)$ are purely real when $\beta,z\in S_{\pm}$, and allow analytic continuation to the whole complex plane, with a power-law bound at infinity, and a branch cut along the positive real axis. Given all the square-root products. we now list the form factors. To further simplify the expressions, we introduce the following dressing functions
$d(\beta)=\frac{\sqrt{\beta}}{\sqrt{N}}\frac{\hat{p}(\beta)}{1+\sqrt{\beta}}\ ,\ \hat{p}(\beta)=p(\beta)(1+i\delta_{\beta,1})\ ,$ (77)
@@
This is the first major result of the work. Furthermore, in terms of the $\hat{\Gamma}_{N}$ function, the ground state overlap has a simple representation
@@
Here we prove the integral representation for $p(\beta)$, $p(z)$ given in Eq. (75) and Eq. (76). We first consider the sum with $\beta\in S_{+}\wedge\beta\neq 1$
@@
Here, the logarithm is in the principal branch and all the square roots are in the $(0,2\pi)$ branch. As an analytic function of $\beta^{\prime}$, there is a branch cut singularity along the positive real axis for the logarithm $\ln\left(1+\frac{\sqrt{\beta^{\prime}}}{\sqrt{\beta}}\right)$, but no other singularities when $\beta^{\prime}$ is away from the positive real axis. As such, one can introduce the following contour representation
@@
Similarly, for the sum ($\beta\in S_{+}\wedge\beta\neq 1$)
@@
which is the desired Eq. (75) for $\cal Z=\beta$. Notice that the extra $\ln(1+\frac{1}{\sqrt{\beta}})$ is due to the $\beta^{\prime}=1$ term in the Eq. (59). The case for ${\cal Z}=z\in S_{-}$ can be established exactly the same way.
@@ 1.4 Scaling limits of the form factors
The integral representations allows to take the scaling limit of the form factors. The crucical quantity is the function $\hat{\Gamma}_{N}({\cal Z})$ defined in Eq. (76). To obtain its scaling limit, one changes the variable as $t\rightarrow e^{-\frac{t}{N}}$, leading to ($\alpha\in R$)
@@
and further seperate the non-zero quantization levels in the scaling limit to the plus ($L$) and minus ($R$) sectors
@@
Then, the dressing function approaches the following limit
@@
Notice that $d(1)=1$ is a constant. In term of the above, the form factors in the scaling limit can be further simplified.
@@
$Z_{-,+;+,+}=Z_{+,-;-,-}=e^{i\frac{\pi}{4}}\ ;\ Z_{-,+;-,-}=-Z_{+,-;+,+}=e^{-i\frac{\pi}{4}}\ ,$
$Z_{0,+,+;+,+}=Z_{0,-,-;-,-}=i\;\ Z_{0,-,-;+,+}=-Z_{0,+,+;-,-}=1\ .$ (109)
To obtain the phase factors, one simply notice that in the scaling limit, the two particle functions in the R sector simplifies as (formulas with in the NS sector are similar)
@@
For the two particle functions involving $\beta=1$, simply notice that in our ordering they all become 1-1. Collecting all the 1-1 and the $e^{\pm\frac{i\pi}{4}}$ factors leads to the final expressions Eq. (109) for the phase factors. The explicit expressions Eq. (107), Eq. (108) and Eq. (109) for the form factors in the scaling limit are the second major result of the paper. They provide explicit formulas for the CFT matrix elements computed recursively in Yurov and Zamolodchikov (1991). Notice that all the form factors in the scaling limit are rational numbers in our convention, due to the overall $e^{\pm\frac{i\pi}{4}}$ in the phase factors.
@@ 2 Derivation of the main results
After introducing all the main results, in this section we derive the results presented in the previous section. The overall strategy is similar to Abraham (1978a, b) [^4] This strategy is essentially equivalent to that of Iorgov and Lisovyy (2011), but allows to avoid the explicit Gaussian wave functions of the ground states. , but the way of solving the matrix equation is different. Rather than using the Wiener-Hopf trick suitable for the infinite-volume limit, our approach is based on the near-Cauchy property of the $W$, $W^{\prime}$ matrices given below in Eq. (128), which is a special property of the critical model and allows explicit expressions at finite $N$.
@@
Since both of the $|\Omega_{R}\rangle$ and $|\Omega_{NS}\rangle$ are gaussian-type ground states of fermionic quadratic forms defined in terms of the same set of fermionic operator, multi-point correlators between the two states are all generated by the overlapping matrices through Wick-contractions Perk et al. (1984). In particular, since the $\sigma^{x}_{0}=\Gamma_{0}$ is also a fermionic operator, correlators with $\sigma$ can also be obtained from Wick contractions. It turns out that it is more convenient to introduce the following matrices
@@
where we have used the mode decompositions in Eq. (34) and Eq. (36). As such, given the overlapping matrices, in the second stage we perform the Wick contraction to obtain the result Eq. (79) of the form factors. Finally, in the third stage, we derive the Eq. (80) for the ground state overlap $|\langle\Omega_{R}|\Omega_{NS}\rangle|$ as an application of the form factors we derived. For notation simplicity, below we will use following notation for the average
@@
where $C(\bar{z})=-i\varphi(z)$. Here we show that, if $W$ is invertible, then the ground state overlap $\langle\Omega_{R}|\Omega_{NS}\rangle$ must be non-vanishing. Assume the opposite, then from Eq. (127), the overlapping matrix with two $c_{\beta}$s must be identical to zero. Then, using the same trick for
@@
and move the $c_{\beta}^{\dagger}$s in $c_{\bar{z}}$ to the left, one can show that all the $\langle\Omega_{R}|c_{\beta_{1}}c_{\beta_{2}}c_{\beta_{3}}c_{\beta_{4}}|\Omega_{NS}\rangle$ are also zero. Recursively [^5] This method can also be used to prove the Wick theorem in the current context, as it leads to a recursive relation for Pfaffians Abraham (1978a)., this implies that all the matrix elements $\langle\Omega_{R}|c_{\beta_{1}}c_{\beta_{2}}c_{\beta_{3}}...c_{\beta_{2n}}|\Omega_{NS}\rangle$ are zero. But this is impossible, because $c_{\beta_{1}}^{\dagger}c^{\dagger}_{\beta_{2}}...c^{\dagger}_{\beta_{n}}|\Omega_{R}\rangle$ is a complete set of orthogonal basis of the full $2^{N}$ dimensional Hilbert space, and vanishing of all the $\langle\Omega_{R}|c_{\beta_{1}}c_{\beta_{2}}c_{\beta_{3}}...c_{\beta_{2n}}|\Omega_{NS}\rangle$ (odd numbers naturally vanish due to the charge condition) would imply that $|\Omega_{NS}\rangle\equiv 0$. We will show below that $W$ is indeed invertible. As such, by dividing the $\langle\Omega_{R}|\Omega_{NS}\rangle$ on the two sides of Eq. (127) one obtains the matrix equation
@@ 2.1 Inversion and summation
Up to now, the discussions are very general and applies to the Ising chain in the whole disordered region, with the $h$-dependent $\varphi$ functions in Eq. (28). At the critical point, however, there is a major simplification: the matrices $W$ and $W^{\prime}$ are nearly Cauchy, up to rank-one perturbations. This allows to obtain explicit product formulas at a generic $N$.
Indeed, using the fact that $\varphi(z)=\sqrt{z}$ when $z\neq 1$, and $\varphi(1)=i$, it is easy to see that one has
@@
Notice that no complex conjugation was performed here, $\langle v|=v^{T}$ means the transpose of $v$. As such, we can introduce the following Cauchy matrices
@@
where $\eta\notin\{-y_{j}\}_{j}$ and $\zeta\notin\{-x_{i}\}_{i}$ are two extra numbers. This formula can be obtained from the standard inversion formula for Cauchy matrices by enlarging the original matrix $(x_{i}+y_{j})^{-1}$ with one extra column and row, with two elements $\eta,\zeta$, and express its inverse at the $\zeta,\eta$ location in two different ways. By taking limits on $\eta$ and $\zeta$, this formula allows to compute all the summations required by this work into products. We first consider the product $HT^{-1}$. This quantity is also required in
@@
We first present the results of $g$. For $\beta,\beta^{\prime}\neq 1$ one has,
@@
$=-\frac{p(\beta)q(\beta^{\prime})}{\sqrt{\beta}+\sqrt{\beta^{\prime}}}-\frac{i-1}{2(1+i)}\bigg(\delta_{\beta,1}-\frac{p(\beta)q(1)}{\sqrt{\beta}+1}\bigg)\bigg(\delta_{\beta^{\prime},1}+\frac{p(1)q(\beta^{\prime})}{\sqrt{\beta^{\prime}}+1}\bigg)$
$=-\frac{\sqrt{\beta^{\prime}}p(\beta)p(\beta^{\prime})(1+\sqrt{\beta}\sqrt{\beta^{\prime}})}{N(1+\sqrt{\beta})(1+\sqrt{\beta^{\prime}})(\sqrt{\beta}+\sqrt{\beta^{\prime}})}\ ,\ \beta,\beta^{\prime}\neq 1\ ,$ (149)
@@
$g_{1,\beta}=-\frac{\sqrt{\beta}p(\beta)}{\sqrt{N}(1+\sqrt{\beta})}\ ,\ g_{\beta,1}=i\frac{p(\beta)}{\sqrt{N}(1+\sqrt{\beta})}\ .$ (150)
And $g_{11}=0$. We now need the following property
$p\left(\bar{\beta}\right)=-i\frac{\sqrt{\beta}-1}{\sqrt{\beta}+1}p(\beta)\ ,$ (151)
which follows from the representation Eq. (75) and the fact that $\hat{\Gamma}_{N}({\cal Z})=\hat{\Gamma}(\bar{\cal Z})$ when ${\cal Z}\in S_{\pm}$. Using this relation, we obtain the desired overlapping matrix in Eq. (112)
@@
As expected, it is manifestly skew-symmetric. Similarly, one obtains the other two overlapping matrices defined in Eq. (113) and Eq. (114). For the $\hat{g}$, one has
$\hat{g}_{z,z^{\prime}}=-\frac{1}{N}\frac{\sqrt{z}}{\sqrt{z}+\sqrt{z^{\prime}}}\frac{1}{p(z)p(z^{\prime})}-\frac{1}{\sqrt{N}}\frac{\sqrt{z}}{(\sqrt{z}-1)(\sqrt{z^{\prime}}-1)p(z)p(z^{\prime})}$

### v4

@@ 1.3 Square-root products and form-factors at finite NN
For the convenience of readers, below we provide a result list for all the computed form-factors. The results at a generic $N$ are given by Eq. (81), and their scaling limits are in Eq. (109), Eq. (110). It turns out that certain square-root products and their relations are crucial in all the stages of the derivation and appear also in the final results, so we decide to interlude for a while, by introducing first all the required square-root products. In our algorithm, they will be naturally generated after inverting certain Cauchy matrices and performing further summations. For $\beta\in S_{+}$, one needs two products
@@
They will be used latter throughout the derivation, in particular, to show a crucial matrix is invertible. On the other hand, for $\beta\neq 1$, no such simple formulas can be derived for the $p$ function. From the perspective of the $S^{2N}_{\pm}$ circles, this is because all the products are restricted to the upper half-plane. But for both $S_{+}$ and $S_{-}$, the $p$ function can still be simplified into a single integral representation. What one needs are the following
@@
where $\cal Z$} can be both $\beta\in S_{+}/\{1\}$ and $z\in S_{-}$. In the expressions, the logarithms are define with the principal branch. The function $\hat{\Gamma}_{N}({\cal Z})$ is purely real when ${\cal Z}\in(S_{+}/\{1\})\cup S_{-}$, and allows analytic continuation to the whole complex plane, with a power-law bound at infinity, and a branch cut along the positive real axis.
Given all the square-root products. we now list the form factors. To further simplify the expressions, we introduce the following dressing functions
$d(\beta)=\frac{\sqrt{\beta}}{\sqrt{N}}\frac{\hat{p}(\beta)}{1+\sqrt{\beta}}\ ,\ \hat{p}(\beta)=p(\beta)(1+i\delta_{\beta,1})\ ,\ d(1)=1\ ,$ (79)
@@
Notice that within $\{\beta_{i}\}$ at most there can be only one $\beta=1$, and all the denominators are well defined. This is the first major result of the work. Furthermore, in terms of the $\hat{\Gamma}_{N}$ function, the ground state overlap has a simple representation
@@
Here we prove the integral representation for $p(\beta)$, $p(z)$ given in Eq. (77) and Eq. (78). We first consider the sum with $\beta\in S_{+}/\{1\}$
@@
Here, the logarithm is in the principal branch and all the square roots are in the $(0,2\pi)$ branch. As an analytic function of $\beta^{\prime}$, there is a branch cut singularity along the positive real axis, but no other singularities when $\beta^{\prime}$ is away from the positive real axis. As such, one can introduce the following contour representation
@@
Similarly, for the sum ($\beta\in S_{+}/\{1\}$)
@@
which is the desired Eq. (77) for ${\cal Z}=\beta\in S_{+}/\{1\}$. Notice that the extra $\ln(1+\frac{1}{\sqrt{\beta}})$ is due to the $\beta^{\prime}=1$ term in the Eq. (61). The case for ${\cal Z}=z\in S_{-}$ can be established exactly the same way.
@@ 1.4 Scaling limits of the form factors
The integral representations allow to take the scaling limit of the form factors. The crucical quantity is the function $\hat{\Gamma}_{N}({\cal Z})$ defined in Eq. (78). To obtain its scaling limit, one changes the variable as $t\rightarrow e^{-\frac{t}{N}}$, leading to ($\alpha\in R_{\neq 0}$)
@@
and further separate the non-zero quantization levels in the scaling limit to the left ($L$) and right ($R$) sectors
@@
Then, the dressing functions approach the following limits
@@
Notice that at the zero momentum, $d(1)=1$ is a constant. In term of the above, the form factors in the scaling limit can be further simplified.
@@
$Z_{-,+;+,+}=Z_{+,-;-,-}=e^{i\frac{\pi}{4}}\ ,\ Z_{-,+;-,-}=-Z_{+,-;+,+}=e^{-i\frac{\pi}{4}}\ ;$
$Z_{0,+,+;+,+}=Z_{0,-,-;-,-}=i\ ,\ Z_{0,-,-;+,+}=-Z_{0,+,+;-,-}=1\ .$ (111)
To obtain the phase factors, one simply notice that in the scaling limit, the two particle functions in the R sector simplify as (formulas with in the NS sector are similar)
@@
For the two particle functions involving $\beta=1$, simply notice that in our ordering they all become 1-1. Collecting all the 1-1 and the $e^{\pm\frac{i\pi}{4}}$ factors leads to the final expressions Eq. (111) for the phase factors. The explicit expressions Eq. (109), Eq. (110) and Eq. (111) for the form factors in the scaling limit are the second major result of the paper. They provide explicit formulas for the CFT matrix elements computed recursively in Yurov and Zamolodchikov (1991), up to the phase conventions. Notice that all the form factors in the scaling limit are rational numbers in our convention, due to the overall $e^{\pm\frac{i\pi}{4}}$ in the phase factors.
Given the matrix elements in the fermionic basis, one can establish the form factor expansions for the multi-point correlators. We will show in Sec. 2.4 that the form factor expansion of the rescaled two-point correlator in the $NS$ ground state can be summed into a product of two Fredholm determinants of the form Eq. (1) given in the abstract, one for the left sector, another for the right sector. Assuming the identity $\det(1+{\cal K})=(1-w)^{-\frac{1}{8}}$ checked up to $w^{30}$ in the small-$w$ expansion, this leads exactly to the standard CFT 2pt on a cylinder. This is another major result of the work.
@@ 2 Derivation of the main results
After introducing all the main results, in this section we derive the results presented in the previous sections. The overall strategy is similar to Abraham (1978a, b) [^4] This strategy is essentially equivalent to that of Iorgov and Lisovyy (2011), but allows to avoid the explicit Gaussian wave functions of the ground states. , but the way of solving the matrix equation is different. Rather than using the Wiener-Hopf trick suitable for the infinite-volume limit, our approach is based on the near-Cauchy property of the $W$, $W^{\prime}$ matrices given below in Eq. (130). This special property of the critical model allows to obtain explicit product formulas at finite $N$ using relatively elementary methods, bypassing the elliptic Cauchy determinant identities Iorgov and Lisovyy (2011) required in more general cases. On the other hand, $W$ and $W^{\prime}$ are Cauchy only up to rank-one perturbations at $\beta=1$, due to the subtle behavior of the critical model at the zero momentum. Not only the square-root function $\varphi(\beta)=\sqrt{\beta}$ for $\beta\in S_{+}/\{1\}$ has a branch cut along the positive real axis, but also, $\varphi(1)=i$ can never be the upper or lower limits of the square-root at $\beta=1$: otherwise the first column of the matrix $\Omega$ in Eq. (23), Eq. (24) would vanish, and $\Omega$ would no-longer be a rotation. This is another special feature of the critical model.
@@
Since both of the $|\Omega_{R}\rangle$ and $|\Omega_{NS}\rangle$ are Gaussian-type ground states of fermionic quadratic forms defined in terms of the same set of fermionic operators, multi-point correlators between the two states are all generated by the overlapping matrices through Wick-contractions Perk et al. (1984). In particular, since the $\sigma^{x}_{0}=\Gamma_{0}$ is also a fermionic operator, matrix elements with $\Gamma_{0}$ can also be obtained from Wick contractions. It turns out that it is more convenient to introduce the following matrices
@@
where we have used the mode decompositions in Eq. (40) and Eq. (42). As such, given the overlapping matrices and the contractions with $\Gamma_{0}$, in the second stage we perform the Wick contraction to obtain the result Eq. (81) of the form factors. Finally, in the third stage, we derive the Eq. (82) for the ground state overlap $|\langle\Omega_{R}|\Omega_{NS}\rangle|$ as an application of the form factors we derived. For notation simplicity, below we will use following notation for the average
@@
where $C(\bar{\cal Z})=-i\varphi({\cal Z})$. Here we show that, if $W$ is invertible, then the ground state overlap $\langle\Omega_{R}|\Omega_{NS}\rangle$ must be non-vanishing. Assume the opposite, then from Eq. (129), the overlapping matrix with two $c_{\beta}$s must be identical to zero. Then, using the same trick for
@@
and move the $c_{\beta}^{\dagger}$s in $c_{\bar{z}}$ to the left, one can show that all the $\langle\Omega_{R}|c_{\beta_{1}}c_{\beta_{2}}c_{\beta_{3}}c_{\beta_{4}}|\Omega_{NS}\rangle$ are also zero. Recursively [^5] This method can also be used to prove the Wick theorem in the current context, as it lead to recursive relations for Pfaffians Abraham (1978a)., this implies that all the matrix elements $\langle\Omega_{R}|c_{\beta_{1}}c_{\beta_{2}}c_{\beta_{3}}...c_{\beta_{2n}}|\Omega_{NS}\rangle$ are zero. But this is impossible, because $c_{\beta_{1}}^{\dagger}c^{\dagger}_{\beta_{2}}...c^{\dagger}_{\beta_{n}}|\Omega_{R}\rangle$ is a complete set of orthogonal basis of the full $2^{N}$ dimensional Hilbert space, and vanishing of all the $\langle\Omega_{R}|c_{\beta_{1}}c_{\beta_{2}}c_{\beta_{3}}...c_{\beta_{2n}}|\Omega_{NS}\rangle$ (odd numbers naturally vanish due to the charge condition) would imply that $|\Omega_{NS}\rangle\equiv 0$. We will show below that $W$ is indeed invertible. As such, by dividing the $\langle\Omega_{R}|\Omega_{NS}\rangle$ on the two sides of Eq. (129) one obtains the matrix equation
@@ 2.1 Inversion and summation
Up to now, the discussions are very general and apply to the Ising chain in the whole disordered region, with the $h$-dependent $\varphi$ functions in Eq. (28). At the critical point, however, there is a major simplification: the matrices $W$ and $W^{\prime}$ are nearly Cauchy, up to rank-one perturbations. This allows to obtain explicit product formulas at a generic $N$ using elementary methods.
Indeed, using the fact that $\varphi({\cal Z})=\sqrt{{\cal Z}}$ when ${\cal Z}\in(S_{+}/\{1\})\cup\ S_{-}$, and $\varphi(1)=i$, it is easy to see that one has
@@
Notice that no complex conjugation was performed here, $\langle v|=v^{T}$ means the transpose of $v$. As such, we can introduce the following $N\times N$ Cauchy matrices ($H$ should not be confused with the Hamiltonian)
@@
where $\eta\notin\{-y_{j}\}_{j}$ and $\zeta\notin\{-x_{i}\}_{i}$ are two extra numbers. This formula can be obtained from the standard inversion formula for Cauchy matrices by enlarging the original Cauchy matrix $(x_{i}+y_{j})^{-1}$ with one extra column and row into a larger Cauchy matrix, with two new elements $\eta,\zeta$, and express its inverse at the $\zeta,\eta$ location in two different ways. By taking limits on $\eta$ and $\zeta$, this formula allows to compute all the summations required by this work into products. We first consider the product $HT^{-1}$. This quantity is also required in
@@
We first present the results of $g$. Combining the Eq. (139), Eq. (141), Eq. (143), Eq. (145) and Eq. (147) one has
@@
$=-\frac{p(\beta)q(\beta^{\prime})}{\sqrt{\beta}+\sqrt{\beta^{\prime}}}-\frac{i-1}{2(1+i)}\bigg(\delta_{\beta,1}-\frac{p(\beta)q(1)}{\sqrt{\beta}+1}\bigg)\bigg(\delta_{\beta^{\prime},1}+\frac{p(1)q(\beta^{\prime})}{\sqrt{\beta^{\prime}}+1}\bigg)\ .$ (151)
We can write the individual components. For both $\beta^{\prime}\neq 1$ and $\beta\neq 1$, one has
$g_{\beta\beta^{\prime}}=-\frac{\sqrt{\beta^{\prime}}p(\beta)p(\beta^{\prime})(1+\sqrt{\beta}\sqrt{\beta^{\prime}})}{N(1+\sqrt{\beta})(1+\sqrt{\beta^{\prime}})(\sqrt{\beta}+\sqrt{\beta^{\prime}})}\ ;\ \beta,\beta^{\prime}\neq 1\ ,$ (152)
@@
$g_{1,\beta}=-\frac{\sqrt{\beta}p(\beta)}{\sqrt{N}(1+\sqrt{\beta})}\ ,\ g_{\beta,1}=i\frac{p(\beta)}{\sqrt{N}(1+\sqrt{\beta})}\ ;\ \beta\neq 1\ .$ (153)
For both $\beta=\beta^{\prime}=1$, from the Eq. (151) and the relation p(1)q(1)=2ip(1)q(1)=-2i, one can compute
$g_{11}=-\frac{p(1)q(1)}{2}+\frac{1-i}{2(1+i)}\left(1-\frac{p(1)q(1)}{2}\right)\left(1+\frac{p(1)q(1)}{2}\right)\equiv 0\ ,$ (154)
as expected. Given $g_{\beta\beta^{\prime}}$, to convert to the $u(\beta,\beta^{\prime})=g_{\beta\bar{\beta}^{\prime}}$, it is helpful to use the following relation for ${\cal Z}\in S_{\pm},\ {\cal Z}\neq 1$
$p\left(\bar{\cal Z}\right)=-i\frac{\sqrt{{\cal Z}}-1}{\sqrt{{\cal Z}}+1}p({\cal Z})\ ,$ (155)
which follows from Eq. (77) and the fact that $\hat{\Gamma}_{N}({\cal Z})=\hat{\Gamma}(\bar{\cal Z})$ when ${\cal Z}\in S_{\pm}$. Using this relation, we obtain the desired overlapping matrix in Eq. (114), where one of the $\beta,\beta^{\prime}$ are not equal to 11
@@
As expected, it is manifestly skew-symmetric. Notice that $u(1,1)=g_{11}=0$. One can also obtain the other two overlapping matrices defined in Eq. (115) and Eq. (116) in a way similar to $u(\beta,\beta^{\prime})$, and we will omit the computation details. For the $\hat{g}$, one has
$\hat{g}_{z,z^{\prime}}=-\frac{1}{N}\frac{\sqrt{z}}{\sqrt{z}+\sqrt{z^{\prime}}}\frac{1}{p(z)p(z^{\prime})}-\frac{1}{N}\frac{\sqrt{z}}{(\sqrt{z}-1)(\sqrt{z^{\prime}}-1)p(z)p(z^{\prime})}$